OpenStudy (anonymous):

if a 30 foot silo cast a 20 foot shadow what is the approximate measure from the tip of the shadow to the top of the silo?

OpenStudy (anonymous):

1) 30^2+20^2=c^2 or c^2=1300 so c=sqrt 1300 or c=about 36 ft. so 36 ft.

OpenStudy (anonymous):

draw the thing first

OpenStudy (anonymous):

Pythagorean theorem 20^2 + 30^2 = x^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):



OpenStudy (anonymous):